OpenStudy (anonymous):

What is the vertex of the function 2x2 + 16x + 21 = 0?

OpenStudy (anonymous):

first coordinate of the vertex is found by the magic formula \(-\frac{b}{2a}\) in your case it is \(-\frac{16}{2\times 2}=-4\) second coordinate is found by replacing \(x\) by \(-4\) to see what you get

OpenStudy (anonymous):

and not to be picky but \[2x^2 + 16x + 21 = 0\] is not a function, it is a quadratic equation \[f(x)=2x^2 + 16x + 21 \] is a function

OpenStudy (anonymous):

and \[f(-4)=2\times (-4)^2+16\times (-4)+21=-11\]so your vertex is \((-4,-11)\)